Calculate the theoretical yield given the balanced equation below and 50.0 g calcium nitrate  

      and 65.0 g potassium phosphate reacting.

3 Ca(NO3)2 (aq)  +  2 K3PO4 (aq)   ->  Ca3(PO4)2 (s)  +  6 KNO3 (aq)


a)Calculate the grams of calcium phosphate formed if all 50.0 g calcium nitrate react.

b)Calculate the grams of calcium phosphate formed if all 65.0 g potassium phosphate react. 

c)   What is the theoretical yield? 

 d) What is the limiting reagent? 


1
Expert's answer
2021-03-15T09:43:32-0400

3 Ca(NO3)2 (aq)  +  2 K3PO4 (aq)   ->  Ca3(PO4)2 (s)  +  6 KNO3 (aq)

3 moles of calcium nitrate = 1 mole of calcium phosphate

3 × 164 g of calcium nitrate = 310g of calcium phosphate

1g calcium nitrate = 310/(3×164) g calcium phosphate

50g calcium nitrate = 310× 50/(3×164) g calcium phosphate

= 31.50g of calcium phosphate


2 moles of potassium phosphate = 1 mole of calcium phosphate

2 × 212g of potassium phosphate = 310 g of calcium phosphate

1g potassium phosphate = 310/(2×212)g calcium phosphate

65g potassium phosphate = 310×65/(2×212)g calcium phosphate

= 47.5 g of calcium phosphate


Limiting reagent is calcium nitrate


Theoretical yield for calcium phosphate = 31.5g

Theoretical yield for potassium nitrate = 6 ×101×50/(3×164)

= 61.59g